# The hereditariness problem for the Černý conjecture

### v1

@@ 1 Introduction
Thus, the analysis of the Černý conjecture can be restricted to three classes of automata: the simple, the radical, and the quasi-simple ones. The latter two classes are described in detail in the sequel. For a preliminary overview of all the classes discussed in this work, see Fig. 1.
The paper is organized as follows. We begin in Section 3 by tackling the problem with a “brute-force” approach, namely by applying the well-known Pin–Frankl algorithm. It turns out that this technique is still insufficient to fully resolve the conjecture. We then move on to study the lattice of congruences of a given automaton. In Section 4 we prove some basic facts about lattices of automaton congruences, with particular focus on the problem of calculating the atoms of such a lattice, which is crucial for computing the radical $\text{Rad}(\mathcal{A})$ of a reset automaton $\mathcal{A}$. We also analyze in more detail the interplay between semisimplicity and the lattice of congruences, obtaining results that will later be used to establish the three-class reduction. In Section 5 we show a Galois connection between this lattice and the lattice of ideals of the transition monoid. This construction not only provides a recursive algorithm to compute the radical $\text{Rad}(\mathcal{A})$ (developted in Section 6), but, more importantly, offers a refined way to tackle the hereditariness problem. In particular, it is fundamental in proving our main results. The remainder of Section 7 is devoted to the study of these classes of automata. The quasi-simple case appears to be the more rigid of the two new classes, and we derive several structural results for it. The central result is Theorem 7.14, which shows that both simple and quasi-simple automata share the property that the transition monoid $\text{M}(\mathcal{A})$ admits a unique maximal ideal lying above $\text{Syn}(\mathcal{A})$. In contrast, for the radical case we were unable to obtain results of comparable depth. Most notably, Proposition 7.6 shows that in a radical automaton the ideal $\text{Syn}(\mathcal{A}/\sigma)$—where $\sigma$ is the minimal non-trivial congruence of $\text{Cong}(\mathcal{A})$—contains a unique non-trivial minimal ideal of $\text{M}(\mathcal{A})/\text{Rad}(\mathcal{A}^{*})$. Finally, Theorem 6.3 is a general result that bounds the index of nilpotency of $\text{Rad}(\mathcal{A}^{*})$ by the height of the congruence lattice $\text{Cong}(\mathcal{A})$, revealing an interesting connection between the combinatorial structure of the automaton $\mathcal{A}$ and the algebraic structure of its transition monoid $M(\mathcal{A})$.
@@ 2 Prerequisites
For a set $A$, we denote $|A|$ the cardinality of $A$. The free monoid on $\Sigma$ is denoted by $\Sigma^{*}$. An ideal $I$ of a monoid $M$ is a subset of $M$ satisfying $MIM\subseteq I$. All semigroup-theoretic notions used in this paper can be found in standard references on semigroup theory, such as [15]. In what follows, an automaton $\mathcal{A}=(Q,\Sigma,\delta)$ will always refer to a deterministic finite automaton (DFA), where $Q$ and $\Sigma$ are finite sets, denoting respectively the set of states and the alphabet, and $\delta:Q\times\Sigma\to Q$ is the transition function. We denote by $|\mathcal{A}|=|Q|$ the number of states of the automaton. Rather than the functional notation, we shall adopt an action notation, writing $q\cdot x$ in place of $\delta(q,x)$. This action extends naturally, first to $\Sigma^{*}$, and then to any subset $F\subseteq Q$, in the obvious way by setting for any $u\in\Sigma^{*}$
$F\cdot u=\{q\cdot u\mid q\in F\}.$
It is a well-known fact that the set
$\text{Syn}(\mathcal{A})=\{u\in\Sigma^{*}:|Q\cdot u|=1\}$
of the reset (or synchronizing) words of $\mathcal{A}$ is a (two-sided) ideal of the monoid $\Sigma^{*}$ that is also regular as a language over $\Sigma$ (see [30]). An automaton is said to be strongly connected if its associated directed graph is strongly connected. We state here the following result, that will be fundamental in what follows:
@@
An important class of automata is given by the simple ones: an automaton $\mathcal{A}$ is said to be simple if $\mathrm{Cong}(\mathcal{A})=\{\Delta_{\mathcal{A}},\nabla_{\mathcal{A}}\}$. A notable infinite class of simple automata is given by the Černý automata (see [30] for their definition and, for instance, [2] for a proof of their simplicity). We now briefly recall the algebraic notation used throughout the paper. Given an automaton $\mathcal{A}$, there exists an epimorphism $\pi:\Sigma^{*}\rightarrow\text{M}(\mathcal{A})$, where $\text{M}(\mathcal{A})$ is the transition monoid associated to $\mathcal{A}$. For convenience, we sometimes drop $\pi$ and identify ideals in $\Sigma^{*}$ with ideals in $\text{M}(\mathcal{A})$ and vice versa; so for instance if we are in the context of $\text{M}(\mathcal{A})$ we will write $\text{Syn}(\mathcal{A})$ instead of $\pi(\text{Syn}(\mathcal{A}))$. It is well-known that $\text{M}(\mathcal{A})$ embeds as a monoid into the ring of matrices $\mathbb{M}_{n}(\mathbb{C})$, where $n$ is the number of states of $\mathcal{A}$ [25, 5]. By slight abuse of notation, we will consider $\text{M}(\mathcal{A})\subseteq\mathbb{M}_{n}(\mathbb{C})$, so that $\pi:\Sigma^{*}\rightarrow\mathbb{M}_{n}(\mathbb{C})$ [2]. A 0-monoid is a monoid equipped with a zero element 0. In such a structure, any non-trivial minimal ideal $I\neq 0$ is called a 0-minimal ideal. Throughout the paper we will often consider the 0-monoid, defined as the Rees quotient $\mathcal{A}^{*}:=\text{M}(\mathcal{A})/\text{Syn}(\mathcal{A})$. For a 0-monoid $M$, a two-sided ideal $I\subseteq M$ is called nilpotent with index of nilpotency $m$ if $I^{m}=0$, where $m$ is the smallest positive integer with this property. It is straightforward to verify that if two ideals $I$ and $J$ satisfy $I^{k}=J^{k}=0$, then $(I\cup J)^{2k}=0$. Hence, the union of nilpotent ideals is itself a nilpotent ideal. This observation allows us to define the largest nilpotent ideal $\text{Rad}(M)$ of $M$, known as the radical of $M$. In the context of automata, the radical ideal of a synchronizing automaton $\mathcal{A}$ is defined as $\text{Rad}(\mathcal{A}^{*})$ (see [2] for an alternative construction via the Wedderburn–Artin theorem). Sometimes we view this ideal inside the free monoid $\Sigma^{*}$, and we set
$\text{Rad}(\mathcal{A}):=\pi^{-1}(\text{Rad}(\mathcal{A}^{*})),$
which is clearly a two-sided ideal of $\Sigma^{*}$. By a slight abuse of notation, we often write $\text{Rad}(\mathcal{A})$ to refer to either $\text{Rad}(\mathcal{A})$ or $\text{Rad}(\mathcal{A}^{*})$, depending on the context. An automaton is called semisimple if $\text{Rad}(\mathcal{A}^{*})=\{0\}$ (equivalently, if $\text{Rad}(\mathcal{A})=\text{Syn}(\mathcal{A})$).
@@
[32, Corollary 4] Let $\mathcal{A}$ be a simple automaton with a letter of defect $1$. Then, $\mathcal{A}$ is synchronizing.
Several algebraic conditions on an automaton guarantee its simplicity. For instance, in [22] it is shown that if the transition monoid $\text{M}(\mathcal{A})/\text{Syn}(\mathcal{A})$ acts irreducibly, then the automaton is simple (called primitive in the terminology of that paper). A further condition implying the simplicity of an automaton is the existence of a subset of letters in the alphabet that generates a submonoid of $\text{M}(\mathcal{A})$ which is a group acting primitivel on the set of states. This topic has been extensively studied; see, for instance, [3].
@@
such that $\varphi(\Sigma^{*}/\text{Syn}(\mathcal{A}))\cong\mathcal{A}^{*}$, allowing us to view $\mathcal{A}^{*}$ as a submonoid of $\mathbb{M}_{n-1}(\mathbb{C})$. We define $\mathcal{R}(\mathcal{A})$ to be the subalgebra of $\mathbb{M}_{n-1}(\mathbb{C})$ generated by $\mathcal{A}^{*}$; we refer to $\mathcal{R}(\mathcal{A})$ as the synchronized $\mathbb{C}$-algebra associated with $\mathcal{A}$. If $\mathcal{A}$ is semisimple—which, as we will see, includes not only the simple case but also the quasi-simple case—then we can apply the Wedderburn–Artin theorem to $\mathcal{R}(\mathcal{A})$, obtaining the decomposition
@@
By minimality, we must have either $\overline{\varphi}_{i}(I)=\mathcal{I}_{i}$ or $\overline{\varphi}_{i}(I)=0$. Since $I\neq\text{Rad}(\mathcal{A})$, it follows that $\overline{\varphi}_{i}(I)=\mathcal{I}_{i}$ for some $i\in\{1,\ldots,k\}$. This suggests that, in order to analyze the Wedderburn–Artin decomposition of a given automaton, it is natural to focus on the study of the 0-minimal ideals of $\text{M}(\mathcal{A})/\text{Rad}(\mathcal{A})$ (or $\text{M}(\mathcal{A})/\text{Syn}(\mathcal{A})$ in the semisimple case).
@@ 3 A standard approach to the hereditariness of the Černý conjecture
In this section, we analyze standard techniques used to address the hereditariness of the Čern’y conjecture. We show that these methods face significant limitations and cannot be extended to a general proof, which motivates the need for a new approach. The challenges discussed here provide the basis for the more algebraic perspective developed in the following sections.
@@
Let $\mathcal{A}$ be a strongly connected synchronizing DFA and $\sigma\in\text{Cong}(\mathcal{A})$ a non-trivial congruence. Assume that $\sigma$ admits a $1$-class or a 22-class: then if $\mathcal{A}/\sigma$ satisfies the Černý conjecture, we have that $\mathcal{A}$ satisfies the Černý conjecture as well.
@@
  • assume that $\sigma$ admits a $2$-class $p]_{\sigma}$ for some $p\in Q$ and no 11-classes. By the same technique used in the previous point, we obtain a word $u\in\text{Syn}(\mathcal{A}/\sigma)$ with $Q\cdot u\subseteq[p]_{\sigma}=\{p,q\}$. By means of the Pin-Frankl Algorithm (see [30]) we have that $\exists u_{0}\in\Sigma^{*}$ such that $|\{p,q\}\cdot u_{0}|=1$ and:
    $|u_{0}|\leq{n\choose 2}=\frac{n(n-1)}{2}=\frac{n^{2}-n}{2}$
    and thus we have that $uu_{0}\in\text{Syn}(\mathcal{A})$ with:
    $|uu_{0}|\leq(n/2-1)^{2}+n/2-1+\frac{n^{2}-n}{2}=\frac{3n^{2}}{4}-\frac{3n}{2}<(n-1)^{2}$
@@
The above case illustrates that we indeed encounter challenges when dealing with a “large” value of $m$. In general, we have the following:
@@
$|u|\leq\frac{n^{2}}{m^{2}}-\frac{n}{m}+\frac{n^{3}-n}{6}-\frac{(n-m)^{3}+3(n-m)^{2}+2(n-m))}{6}.$
@@
Let $u_{0}\in\text{Syn}(\mathcal{A}/\sigma)$ satisfying Černý and $u_{1}\in\Sigma^{*}$ such that $Q\cdot u_{0}u_{1}\subseteq[p]_{\sigma}$ with $|[p]_{\sigma}|=m$. With the same technique adopted in the second case of the previous proof, again by the Pin-Frankl algorithm we have that $\exists v\in\Sigma^{*}$ such that $|[p]_{\sigma}\cdot v|=1$ and:
$\sum_{k=2}^{m}{n-k+2\choose 2}=\frac{n^{3}-n}{6}-\frac{(n-m)^{3}+3(n-m)^{2}+2(n-m)}{6}$

### v2

@@ 1 Introduction
Thus, the analysis of the Černý conjecture can be restricted to three classes of automata: the simple, the radical, and the quasi-simple ones. The latter two classes are described in detail in the sequel. It is worth noting that, when automata are viewed as unary algebras, these two classes are strictly contained in the class of subdirectly irreducible algebras; see [8, Theorem 8.4]. For a preliminary overview of all the classes discussed in this work, see Fig. 1.
The paper is organized as follows. We begin in Section 3 by tackling the problem with a “brute-force” approach, namely by applying the well-known Pin–Frankl algorithm. It turns out that this technique is still insufficient to fully resolve the conjecture. We then move on to study the lattice of congruences of a given automaton. In Section 4 we recall some basic facts about lattices of automaton congruences, with particular focus on the problem of calculating the atoms of such a lattice, which is crucial for computing the radical $\text{Rad}(\mathcal{A})$ of a reset automaton $\mathcal{A}$. We also analyze in more detail the interplay between semisimplicity and the lattice of congruences, obtaining results that will later be used to establish the three-class reduction. In Section 5 we show a Galois connection between this lattice and the lattice of ideals of the transition monoid. This construction not only provides a recursive algorithm to compute the radical $\text{Rad}(\mathcal{A})$ (developted in Section 6), but, more importantly, offers a refined way to tackle the hereditariness problem. In particular, it is fundamental in proving our main results. The remainder of Section 7 is devoted to the study of these classes of automata. The quasi-simple case appears to be the more rigid of the two new classes, and we derive several structural results for it. The central result is Theorem 7.14, which shows that both simple and quasi-simple automata share the property that the transition monoid $\text{M}(\mathcal{A})$ admits a unique maximal ideal lying above $\text{Syn}(\mathcal{A})$. In contrast, for the radical case we were unable to obtain results of comparable depth. Most notably, Proposition 7.6 shows that in a radical automaton the ideal $\text{Syn}(\mathcal{A}/\sigma)$, where $\sigma$ is the minimal non-trivial congruence of $\text{Cong}(\mathcal{A})$ (also called monolith of $\mathcal{A}$), contains a unique non-trivial minimal ideal of $\text{M}(\mathcal{A})/\text{Rad}(\mathcal{A}^{\star})$. Finally, Theorem 6.4 is a general result that bounds the index of nilpotency of $\text{Rad}(\mathcal{A}^{\star})$ by the height of the congruence lattice $\text{Cong}(\mathcal{A})$, revealing an interesting connection between the combinatorial structure of the automaton $\mathcal{A}$ and the algebraic structure of its transition monoid $\text{M}(\mathcal{A})$.
@@ 2 Prerequisites
For a set $A$, we denote $|A|$ the cardinality of $A$. The free monoid on $\Sigma$ is denoted by $\Sigma^{*}$. An ideal $I$ of a monoid $M$ is a subset of $M$ satisfying $MIM\subseteq I$. All semigroup-theoretic notions used in this paper can be found in standard references on semigroup theory, such as [18]. In what follows, an automaton $\mathcal{A}=(Q,\Sigma,\delta)$ refers to a deterministic finite automaton (DFA), where $Q$ and $\Sigma$ are finite sets denoting the set of states and the alphabet, respectively, and $\delta:Q\times\Sigma\to Q$ is the transition function. We denote by $|\mathcal{A}|=|Q|$ the number of states of the automaton. Rather than the functional notation, we adopt an action notation, writing $q\cdot x$ in place of $\delta(q,x)$. This action extends naturally to the free monoid $\Sigma^{*}$, and then to any subset $F\subseteq Q$, by setting
$F\cdot u=\{q\cdot u\mid q\in u\}$
for any $u\in\Sigma^{*}$. The transition monoid of $\mathcal{A}$, denoted by $\text{M}(\mathcal{A})$, is the quotient of $\Sigma^{*}$ by the congruence $\equiv_{\mathcal{A}}$ defined by $u\equiv_{\mathcal{A}}v$ if and only if $q\cdot u=q\cdot v$ for all $q\in Q$ (i.e., the kernel of the action). Henceforth, we will consider automata that are synchronizing. It is a well-known fact that the set
$\mathrm{Syn}(\mathcal{A})=\{u\in\Sigma^{*}:|Q\cdot u|=1\}$
of the reset (or synchronizing) words of $\mathcal{A}$ is a two-sided ideal of the free monoid $\Sigma^{*}$. Furthermore, $\mathrm{Syn}(\mathcal{A})$ is a regular language over $\Sigma$ (see [33]). An element $u\in\text{Syn}(\mathcal{A})$ satisfying the Černý bound, i.e. such that $|u|\leq(n-1)^{2}$ is said to be a Černý reset word. An automaton is said to be strongly connected if its associated directed graph is strongly connected. We state here the following result, that will be fundamental in what follows:
@@
An important class of automata is given by the simple ones: an automaton $\mathcal{A}$ is said to be simple if $\mathrm{Cong}(\mathcal{A})=\{\Delta_{\mathcal{A}},\nabla_{\mathcal{A}}\}$. A notable infinite class of simple automata is given by the Černý automata (see [33] for their definition and, for instance, [2] for a proof of their simplicity). We now briefly recall the algebraic notation used throughout the paper. Given an automaton $\mathcal{A}$, there exists an epimorphism $\pi:\Sigma^{*}\rightarrow\text{M}(\mathcal{A})$, where $\text{M}(\mathcal{A})$ is the transition monoid associated to $\mathcal{A}$. For convenience, we sometimes drop $\pi$ and identify ideals (and elements) in $\Sigma^{*}$ with ideals (and elements) in $\text{M}(\mathcal{A})$ and vice versa; so for instance if we are in the context of $\text{M}(\mathcal{A})$ we will write $\text{Syn}(\mathcal{A})$ instead of $\pi(\text{Syn}(\mathcal{A}))$. It is well-known that $\text{M}(\mathcal{A})$ embeds as a monoid into the ring of matrices $\mathbb{M}_{n}(\mathbb{C})$, where $n$ is the number of states of $\mathcal{A}$ [28, 5]. By slight abuse of notation, we will consider $\text{M}(\mathcal{A})\subseteq\mathbb{M}_{n}(\mathbb{C})$, so that $\pi:\Sigma^{*}\rightarrow\mathbb{M}_{n}(\mathbb{C})$ [2]. A 0-monoid is a monoid equipped with a zero element 0. In such a structure, any non-trivial minimal ideal $I\neq 0$ is called a 0-minimal ideal. Throughout the paper we will often consider the 0-monoid, defined as the Rees quotient $\mathcal{A}^{\star}:=\text{M}(\mathcal{A})/\text{Syn}(\mathcal{A})$, and we denote by $\theta:\text{M}(\mathcal{A})\to\mathcal{A}^{\star}$ the corresponding Rees morphism. For a 0-monoid $M$, a two-sided ideal $I\subseteq M$ is called nilpotent with index of nilpotency $m$ if $I^{m}=0$, where $m$ is the smallest positive integer with this property. It is straightforward to verify that if two ideals $I$ and $J$ satisfy $I^{k}=J^{k}=0$, then $(I\cup J)^{2k}=0$. Hence, the union of nilpotent ideals is itself a nilpotent ideal. This observation allows us to define the largest nilpotent ideal $\text{Rad}(M)$ of $M$, known as the radical of $M$. In the context of automata, the radical ideal of a synchronizing automaton $\mathcal{A}$ is defined as $\text{Rad}(\mathcal{A}^{\star})$ (see [2] for an alternative construction via the Wedderburn–Artin theorem). We will need to see this ideal in $\text{M}(\mathcal{A})$ and thus we set:
$\text{Rad}(\mathcal{A}):=\theta^{-1}(\text{Rad}(\mathcal{A}^{\star}))\subseteq\text{M}(\mathcal{A}).$
By a slight abuse of notation, we write $\text{Rad}(\mathcal{A})$ to refer to either $\text{Rad}(\mathcal{A})$ or $\pi^{-1}(\text{Rad}(\mathcal{A}))\subseteq\Sigma^{*}$, depending on the context. An automaton is called semisimple if $\text{Rad}(\mathcal{A}^{\star})=\{0\}$ (equivalently, if $\text{Rad}(\mathcal{A})=\text{Syn}(\mathcal{A})$).
@@
[35, Corollary 4] Let $\mathcal{A}$ be a simple automaton with a letter of defect $1$. Then, $\mathcal{A}$ is synchronizing.
Several algebraic conditions on an automaton guarantee its simplicity. For instance, in [25] it is shown that if the transition monoid $\text{M}(\mathcal{A})/\text{Syn}(\mathcal{A})$ acts irreducibly, then the automaton is simple (called primitive in the terminology of that paper). A further condition implying the simplicity of an automaton is the existence of a subset of letters in the alphabet that generates a submonoid of $\text{M}(\mathcal{A})$ which is a group acting primitively on the set of states. This topic has been extensively studied; see, for instance, [3].
@@
such that $\varphi(\Sigma^{*}/\text{Syn}(\mathcal{A}))\cong\mathcal{A}^{\star}$, allowing us to view $\mathcal{A}^{\star}$ as a submonoid of $\mathbb{M}_{n-1}(\mathbb{C})$. We define $\mathcal{R}(\mathcal{A})$ to be the subalgebra of $\mathbb{M}_{n-1}(\mathbb{C})$ generated by $\mathcal{A}^{\star}$; we refer to $\mathcal{R}(\mathcal{A})$ as the synchronized $\mathbb{C}$-algebra associated with $\mathcal{A}$. If $\mathcal{A}$ is semisimple—which, as we will see, includes not only the simple case but also the quasi-simple case—then we can apply the Wedderburn–Artin theorem to $\mathcal{R}(\mathcal{A})$, obtaining the decomposition
@@
By minimality, we must have either $\overline{\varphi}_{i}(I)=\mathcal{I}_{i}$ or $\overline{\varphi}_{i}(I)=0$. Since $I\neq\text{Rad}(\mathcal{A})$, it follows that $\overline{\varphi}_{i}(I)=\mathcal{I}_{i}$ for some $i\in\{1,\ldots,k\}$. This suggests that, in order to analyze the Wedderburn–Artin decomposition of a given automaton, it is natural to focus on the study of the 0-minimal ideals of $\text{M}(\mathcal{A})/\text{Rad}(\mathcal{A})$ (or $\mathcal{A}^{\star}$ in the semisimple case).
@@ 3 A standard approach to the hereditariness of the Černý conjecture
In this section, we analyze standard techniques used to address the hereditariness of the Černý conjecture. We show that these methods face significant limitations and cannot be extended to a general proof, which motivates the need for a new approach. The challenges discussed here provide the basis for the more algebraic perspective developed in the following sections.
@@
Let $\mathcal{A}$ be a strongly connected synchronizing DFA and $\sigma\in\text{Cong}(\mathcal{A})$ a non-trivial congruence. Assume that $\sigma$ admits a 1-class or a 2-class: then if $\mathcal{A}/\sigma$ satisfies the Černý conjecture, we have that $\mathcal{A}$ satisfies the Černý conjecture as well.
@@
  • assume that $\sigma$ admits a 2-class $p]_{\sigma}$ for some $p\in Q$ and no 1-classes. By the same technique used in the previous point, we obtain a word $v\in\text{Syn}(\mathcal{A}/\sigma)$ with $Q\cdot v\subseteq[p]_{\sigma}=\{p,q\}$. By means of the Pin-Frankl Algorithm (see [33]) we have that $\exists v_{0}\in\Sigma^{*}$ such that $|\{p,q\}\cdot v_{0}|=1$ and:
    $|v_{0}|\leq{n\choose 2}=\frac{n(n-1)}{2}=\frac{n^{2}-n}{2}$
    and thus we have that $vv_{0}\in\text{Syn}(\mathcal{A})$ with:
    $|vv_{0}|\leq(n/2-1)^{2}+n/2-1+\frac{n^{2}-n}{2}=\frac{3n^{2}}{4}-n\leq(n-1)^{2},\mbox{ for }n\geq 2$
@@
The above case illustrates that we indeed encounter challenges when dealing with a “large” value of the cardinality of the smallest equivalence class of $\sigma$. In general, we have the following:
@@
$|u|\leq\frac{n^{2}}{m^{2}}-\frac{n}{m}+\frac{n^{3}-n}{6}-\frac{(n-m)^{3}+3(n-m)^{2}+2(n-m)}{6}.$
@@
Let $u_{0}\in\text{Syn}(\mathcal{A}/\sigma)$ be a Černý reset word for $\mathcal{A}/\sigma$ and $u_{1}\in\Sigma^{*}$ such that $Q\cdot u_{0}u_{1}\subseteq[p]_{\sigma}$ with $|[p]_{\sigma}|=m$. With the same technique adopted in the second case of the previous proof, again by the Pin-Frankl algorithm we have that $\exists v\in\Sigma^{*}$ such that $|[p]_{\sigma}\cdot v|=1$ and:
$|v|\leq\sum_{k=2}^{m}{n-k+2\choose 2}=\frac{n^{3}-n}{6}-\frac{(n-m)^{3}+3(n-m)^{2}+2(n-m)}{6}$